# CMT: A Cascade MAR with Topology Predictor for Multimodal Conditional CAD Generation

### v1

@@ Abstract
While accurate and user-friendly Computer-Aided Design (CAD) is crucial for industrial design and manufacturing, existing methods still struggle to achieve this due to their over-simplified representations or architectures incapable of supporting multimodal design requirements. In this paper, we attempt to tackle this problem from both methods and datasets aspects. First, we propose a cascade MAR [25] with topology predictor (CMT), the first multimodal framework for CAD generation based on Boundary Representation (B-Rep). Specifically, the cascade MAR can effectively capture the “edge-counters-surface” priors that are essential in B-Reps, while the topology predictor directly estimates topology in B-Reps from the compact tokens in MAR. Second, to facilitate large-scale training, we develop a large-scale multimodal CAD dataset, mmABC, which includes over 1.3 million B-Rep models with multimodal annotations, including point clouds, text descriptions, and multi-view images. Extensive experiments show the superior of CMT in both conditional and unconditional CAD generation tasks. For example, we improve Coverage and Valid ratio by +10.68% and +10.3%, respectively, compared to state-of-the-art methods on ABC [21] in unconditional generation. CMT also improves +4.01 Chamfer on image conditioned CAD generation on mmABC. The dataset, code and pretrained network shall be released.
@@ 2.2 CAD Datasets
Figure 2: The pipeline of CMT. CMT consists of continuous tokenization of B-Rep, a unified multimodal condition encoder, a cascade autoregressive generation network, and a topology predictor. Given a CAD model $B$ and its corresponding multimodal inputs $C$, (a) the continuous tokenization process in CMT converts the B-Rep into continuous surface tokens $S$ and edge tokens $E$ using Surface VAE [20] and Edge VAE respectively. (b) the users’ input $C$ is encoded into a fixed-length condition embedding $Z$ in the unified multimodal condition encoder. (c) The edge sequence is randomly masked, and the edge MAR is trained to reconstruct it from learnable masked tokens $M_{e}$ based on visible edge tokens $E_{v}$ and the condition embedding $Z$. The generated varying‑length edge sequence, along with the condition embedding, is further encoded into a fixed-length edge embedding $Q$, which are concatenated with the visible surface tokens $S_{v}$ and masked tokens $M_{s}$ to reconstruct masked surface tokens. (d) a cross-attention layer [39] is utilized to predict the adjacent matrix $\hat{R}$ between the generated edge tokens $\hat{E}$ and the generated surface tokens ${\hat{E}}$. The overall loss comprises the topology error $L_{topo}$ from the adjacency matrix prediction, as well as the token generation loss of edges $L_{edge}$ and surfaces $L_{surf}$.
@@ 4.4 Topology Predictor
In order to reconstruct the final B-Rep, we aim to obtain the topology relationship $R\in\mathbb{R}^{N_{e}\times N_{s}}$ between the generated edge tokens $\hat{E}$ and surface tokens $\hat{S}$ through Topology Predictor. Thanks to the precise generation of edge tokens and surface tokens, we directly predict topology relationship $\hat{R}$ between edges and surfaces with a simple cross-attention layer [39] $\mathcal{G}_{t}$ i.e., $A=\mathcal{G}_{t}(\hat{E},\hat{S})$. A threshold $\tau$ is set to determine whether a relationship exists. When the score in the generated matrix $A\in\mathbb{R}^{N_{e}\times N_{s}}$ exceeds the threshold $\tau=0.5$, it is considered that there is an adjacency relationship between the edge and the surface.

### v2

@@ Abstract
While accurate and user-friendly Computer-Aided Design (CAD) is crucial for industrial design and manufacturing, existing methods still struggle to achieve this due to their over-simplified representations or architectures incapable of supporting multimodal design requirements. In this paper, we attempt to tackle this problem from both methods and datasets aspects. First, we propose a cascade MAR [25] with topology predictor (CMT), the first multimodal framework for CAD generation based on Boundary Representation (B-Rep). Specifically, the cascade MAR can effectively capture the “edge-counters-surface” priors that are essential in B-Reps, while the topology predictor directly estimates topology in B-Reps from the compact tokens in MAR. Second, to facilitate large-scale training, we develop a large-scale multimodal CAD dataset, mmABC, which includes over 1.3 million B-Rep models with multimodal annotations, including point clouds, text descriptions, and multi-view images. Extensive experiments show the superior of CMT in both conditional and unconditional CAD generation tasks. For example, we improve Coverage and Valid ratio by +10.68% and +10.3%, respectively, compared to state-of-the-art methods on ABC [21] in unconditional generation. CMT also improves +4.01 Chamfer on image conditioned CAD generation on mmABC.
@@ 2.2 CAD Datasets
Figure 2: The pipeline of CMT. CMT consists of continuous tokenization of B-Rep, a unified multimodal condition encoder, a cascade autoregressive generation network, and a topology predictor. Given a CAD model $B$ and its corresponding multimodal inputs $C$, (a) the continuous tokenization process in CMT converts the B-Rep into continuous surface tokens $S$ and edge tokens $E$ using Surface VAE [20] and Edge VAE respectively. (b) the users’ input $C$ is encoded into a fixed-length condition embedding $Z$ in the unified multimodal condition encoder. (c) The edge sequence is randomly masked, and the edge MAR is trained to reconstruct it from learnable masked tokens $M_{e}$ based on visible edge tokens $E_{v}$ and the condition embedding $Z$. The generated varying‑length edge sequence, along with the condition embedding, is further encoded into a fixed-length edge embedding $Q$, which are concatenated with the visible surface tokens $S_{v}$ and masked tokens $M_{s}$ to reconstruct masked surface tokens. (d) a cross-attention layer [39] is utilized to predict the adjacent matrix $\hat{R}$ between the generated edge tokens $\hat{E}$ and the generated surface tokens $^{\hat{E}}$. The overall loss comprises the topology error $L_{topo}$ from the adjacency matrix prediction, as well as the token generation loss of edges $L_{edge}$ and surfaces $L_{surf}$.
@@ 4.4 Topology Predictor
[…]
@@ 6 Conclusion
Acknowledgements
This work was supported by the National Key R&D Program of China(2022ZD0160201), Shanghai Artificial Intelligence Laboratory, the JC STEM Lab of AI for Science and Engineering, funded by The Hong Kong Jockey Club Charities Trust, the Research Grants Council of Hong Kong (Project No. CUHK14213224).